OpenStudy (jojo12g):

Divide. –7.2 ÷ –0.9 A. –8 B. –0.8 C. 0.8 D. 8

alones (alones):

8

OpenStudy (anonymous):

change the negative numbers to positive ones then do the equation = 8

OpenStudy (jojo12g):

thanks guys

OpenStudy (welshfella):

-7.2 / -0.9 = 7.2 / 0.9 = 72 / 9 = 8